OpenStudy (anonymous):

I've actually come across a question in my review packet I don't understand how to do, it's show that the function F(x) = ∫x4x1/tdt is constant on the interval 0 to positive infinity.

OpenStudy (anonymous):

\[\int\limits_{x}^{4x} 1/t\]

jimthompson5910 (jim_thompson5910):

I think it wants you to use the fundamental theorem of calculus to find F ' (x) and show that F ' (x) is 0 When F ' (x) = 0, the function F(x) is constant

OpenStudy (anonymous):

I realized that after I posted it, completely forgot about the fundamental theorem, it comes out to ln (4)

jimthompson5910 (jim_thompson5910):

I meant this part of the FTC If \[\Large g(x) = \int_{m(x)}^{n(x)}f(t)dt\] then \[\Large g \ '(x) = f \ '(n(x))*n \ '(x) - f \ '(m(x))*m \ '(x)\] I'm using the chain rule as well

OpenStudy (anonymous):

I think I understand that problem, do you think you could help me with another one?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Find \[\int\limits_{2}^{x^3} \ln(x^2) dx\]

jimthompson5910 (jim_thompson5910):

is that a typo? you used x in two different spots

OpenStudy (anonymous):

That's how my packet has it i'm just as confused as you are, it should be t or something right?

jimthompson5910 (jim_thompson5910):

yeah it should be ln(t^2)dt

OpenStudy (anonymous):

ok so assuming that, how would I solve this?

jimthompson5910 (jim_thompson5910):

does it want you to use the FTC again?

jimthompson5910 (jim_thompson5910):

oh wait, nvm I see how to do this

jimthompson5910 (jim_thompson5910):

ln(t^2) = 2*ln(t) using one of the log rules

jimthompson5910 (jim_thompson5910):

you can then use integration by parts to integrate ln(t)

OpenStudy (anonymous):

This is under the fundamental theorem of calculus section of my packet, how would I use that to solve this?

jimthompson5910 (jim_thompson5910):

what are the full instructions? if possible, take a screenshot or take a pic of the full problem. also, have you learned about integration by parts yet?

OpenStudy (anonymous):

I have learned integration of parts, It only says find (equation) but is under the section of fundamental theorem of calculus, so I assume that I need to use the theorem to solve it,

jimthompson5910 (jim_thompson5910):